Jan 12, 2017 Calorimeter

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 10:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run preparations
  4. FCAL update
  5. BCAL update
    • LED monitoring (Ahmed)
  6. Calibrations
    • BCAL Gain calibration (Will)
  7. Reconstruction and Simulation
  8. NIM articles
  9. Any other business


Attendees: Adesh, Marc, Elton, Jon (JLab); Zisis, Ahmed, Andrei (UofR); Matt (IU); Christina, George (Athens); Will (CMU)

  1. Announcements
    • First meeting of the year!
    • Spring run starts in about 2 weeks.
    • Update on Lustre system: [info from Mark Ito]: Five disk servers (out of about 50-100) failed on Dec 31. Two have been recovered. Three remain down and Intel engineers are working to get them back. Files are spread over the entire disk system, so the impact is proportionally larger than might be expected. The files themselves are thought to be ok so fixing the servers should restore the files. Unfortunately the fixes are labor intensive and will possibly take weeks.
  2. Action Items
  3. Run preparations
    • HCO checklist will be downgraded to 'not ready' in preparation for new run.
  4. BCAL update
    1. LED monitoring (Ahmed)
      • Checking 2016 fall run 21844. Does not look stable.
      • Lustre system has been down and access to files difficult.
      • Will report at the next meeting.
      • Monitoring report GlueX-doc-3191 has been posted on the portal.
    2. Led noise (Marc)
      • During the fall run, four channels (two upstream and two downstream) exhibited increased noise. The only obvious correlation is that they view a particular LED. Will investigate behavior before the next run.
  5. Calibrations
    • BCAL Gain calibration (Will)
    • pi0 skim files for Bcal and Fcal are corrupted (new format problem?) Sean has been informed and will redo the skim. Will also tried running the eviowriter, but gets the same error. New gain calibrations will be obtained as soon as new files are available.
  6. Reconstruction and Simulation
    1. Negative BCAL TDC times in simulation? See e-mail from Richard
      • Richard noticed in Geant4 that mcsmear is cutting off TDC times for negative values. Also the TDC and ADC time offsets seem to be quite different. This is likely a bug of some kind.
      • Will and Sean will try to understand the problem (hopefully fix it).
    2. George: Working on pi0 Dalitz decays and making progress. However Lustre problems have delayed completion. Started looking at MC to optimize cuts.
  7. FCAL update
    1. Humidity sensor lowered to 2% (four measurements in the Fcal dark box) because they were getting close to the updated threshold of 5%. The environment should be back to normal in the hall.
    2. Base repair (Jon)
      • Chris ill continue repairing bases at about 20/week. Pair spectrometer might start taking up more of his time.
      • At the moment we have about 57 spares.
      • Six bases have failed in the past month
      • Continue to monitor the spare pool in the lab. About 3 have failed out of 60 repaired. These three will be send to Paul at IU for diagnostics. In particular we would like to know if fixes and/or technique is insufficient to complete necessary repairs.
      • Fernando continues to take a look at the design and has some thoughts on the use of a replacement transistor to alleviate the failures.
    3. Tracking spread sheet (Matt)
      • Suggests keeping a daily log with a) detector status b) HV status c) number of failed bases on Fcal, d) number of bases on/off in test stand e) HV status of test bases. Elton: Suggest to record temperature and humidity at the same time.
    4. Calibrations (Adesh)
      • Will start with check of the pedestals.
      • Need skim files, but calibration can be completed in a few days once they are available.
      • Elton: Asked about Bcal pedestals. Marc: We subtract event-by-event, which is automatic. However, the new FADC format includes more information. Pedestal events might also be added to data stream. We should check and incorporate into monitoring/data checks.
    5. FADC problem in Fcal 10
      • Fadc card had problem running at high rate, but otherwise seemed fine. Sasha identified the problem. Ben or Ed are woking on a way to checking that type of problem routinely.
  8. NIM articles
    • Fcal and Bcal NIM articles are up at the top of our agendas and will be followed
    • Elton suggested using the private git repository for them.
  9. Any other business